Grade 12IIT JEE Entrance Exam

7.* By which of the following methods, H2O2 can be synthesised ?
(A) Addition of ice cold H2SO4 on BaO

prakhar kumar
8 Years agoGrade 12

1 Answer

ramteja
ApprovedApproved Tutor Answer8 Years ago
.* F2 + dil NaOH  A + NaF + H2O ; F2 + conc. NaOH  B + NaF + H2O A & B are : (A) A = OF2 (B) A = O2 (C) B = O2 (D) B = OF2 7.* By which of the following methods, H2O2 can be synthesised ? (A) Addition of ice cold H2SO4 on BaO2 (B) Addition of ice cold H2SO4 on PbO2 (C) Aerial oxidation of 2-ethyl anthraquinol (D) Electrolysis of (NH4 ) 2SO4 at a high current density